# Design Principles of Zero-Shot Self-Supervised Unknown Emitter Detectors

### v1

@@ III-B Same Messages Scenario (SM)
$\begin{split}&x_{i}(t)-x^{\prime}_{i}(t)\approx 0\\ &x_{i}(t)-x_{j}(t)=r_{emitter,i}(t)-r_{emitter,j}(t)\neq 0,\end{split}$ (4)
which makes the self-supervised task easier because of linearly separable classes.
III-C Different Messages Scenario (DM)
However, in more realistic use cases whem we don’t know the exact protocol, we only have access to some random part of WiFi frame, which means that clean signal $x_{symbols}$ is random. The signal can now be expressed as:
@@ III-C Different Messages Scenario (DM)
where $\hat{x}_{symbols}(t)$ is random. Thus, the variation inside emitter becomes the same as differences between emitters because the emiiter specific part is a small perturbation comparing to the clean symbol as stated in Eq. 2:
@@
Figure 2: Yellow dots represent an arbitrary signal from the first emitter from the WiSig Dataset [hanna_wisig_2022], and black dots stand for the arbitrary signal from the second Emitter. The signals carry the same message and differ inly by environmental noise and hardware imperfections
Figure 3: SEI workflow of the proposed approach
@@ III-D Problem Statement
In this work, we formulate the problem of unknown emitter identification as a zero-shot learning task within a self-supervised learning setting. We build an approach that extracts the RF Fingerprint part from signals, Eq. LABEL:eq:rf_distortion, grouping the emitters in the features space.
The objective is to learn a composite function
$\Phi:X_{\text{in}}\rightarrow L,\quad\text{where }L=\{\text{known}\>(0),\text{unknown}\>(1)\},$ (7)
@@
The composite function $\Phi(X_{\text{in}})$ from Eq. (7) can be decomposed into two modular sub-functions as per Eq. (8): the embedding function $F$, which transforms raw RF data into an embedding in a discriminative feature space, and the detection function $M$, which determines whether an embedded sample belongs to a known or an unknown emitter.
@@ III-D1 Embedding Function
Let $F$ be the embedding function trained in a self-supervised manner to project input signals into a latent space:
$F:X_{\text{in}}\rightarrow\mathbb{R}^{d}$ (9)
where $X_{\text{in}}$ denotes the input signal domain, and $\mathbb{R}^{d}$ is the latent feature space of dimension $d$. The output embeddings should cluster tightly for samples from the same emitter while remaining well-separated for different emitters.
@@ III-D2 Detection Function
The detection function $M$ operates over the embedded space to determine whether a given sample is associated with a known or unknown emitter:
$M:\mathbb{R}^{d}\rightarrow L$ (10)
IV Design of the SEI workflow
To tackle the problem identified in Section III-D, we propose a new approach according to the workflow depicted in Figure 3. Firstly, the receiver gets signals from the transmitting RF emitters in the Raw I/Q form. Then, the signal processing step normalizes the data and transforms it to the desired Data Modality. After that, the Feature Learning module $F$, Eq. 9, trained by one of the Self-Supervised Learning approaches produces the embeddings of the input samples. Next, the Decision Making Module (DMM) gives a classification decision based on the extracted features. Finally, we also employ different complementary metrics to evaluate the performance of the proposed workflow. The following subsections describe each of these steps in details.
@@ IV-A Transmitting devices
In this study devices transmit signals following the two scenarios described in Section III-D: the Same Messages (Section III-B) and Different Messages (Section III-C). The first scenario in which all emitters transmit the same content all the time is considered in case when one has access to preamble. The second scenario with random content is more realistic but also more challenging. The scenario affects the best design choices of Unknown Emitter Self-Supervised Zero-Shot detectors.
@@
IV-B Data Modality
When the RF signal comes to receiver, it is descretizded with the same time step: $x[i]=x(i\delta t)$, where $x[i]$ is a complex value denoting I and Q components. The way how it is processed to be fed to features learning module $F$ is called a Data Modality - an important of the design space. In the current work, we use two types of data modalities: Raw I/Q and the proposed 2D constellation. The Raw I/Q is the convenient choice for OFDM protocol data; each observation $x[i]=x[i]^{I}+jx[i]^{Q}\in\mathbb{C}^{N}$ is represented as a two-channel input:
$X_{\text{in}}^{0i}=x^{I}[i],\quad X_{\text{in}}^{1i}=x^{Q}[i]$ (11)
In this study, we argue that Raw I/Q representation produces poor quality under the realistic scenario assuming different transmitting messages because of the prevailing of semantic information $x_{symbols}(t)$ over the fingerprint information $r_{emittter}(t)$ (Eq. 5).
To address this issue, one can transform the data in time-invariant manner:
@@ IV-B Data Modality
where $T$ is a transformation operation over raw I/Q data, and $\sigma$ is a permutation operation over the time axis. The $T(X_{\text{in}})$ tensor does not contain any semantic information, because all messages have the same representation, assuming that transmitted signals are random. The simplest way to provide the time-invariant property is to represent raw I/Q samples as 2D Constellation. Firstly, the data is scaled to [1,1][-1,1] range:
@@
Secondly, the procedure counting I/Q observations over cells created by the grid in a 2D plane (Figure 4) is applied:
$\begin{split}T(X_{\text{in}})^{i,j}=\#\{k:\frac{X_{\text{in norm}}^{0k}+1}{2}\in[i\epsilon,(i+1)\epsilon],\\ \frac{X_{\text{in norm}}^{1k}+1}{2}\in[j\epsilon,(j+1)\epsilon]\}/N\\ \text{where }i,j\in[0,K-1],\epsilon=\frac{1}{K}\end{split},$ (14)
where $K$ is the grid size and $N$ is the number of I/Q observations.
@@ IV-C Feature Learning Module
The feature learning module $F$ is a core part of Unknown Emitter Self-Supervised Zero-shot detectors. One of the most common architectures of feature learning module are Convolutional Neural Netowrks (CNNs), which are used in SOTA approaches [liu2023overcoming, hao2023contrastive]. We implement 1D and 2D CNN architectures for raw I/Q and 2D-Constellation data modalities, respectively. The 1D CNN two channels network is adopted from [o2018over] and consists of four blocks, each containing convolution, batch normalization and max pooling operations. 2D CNN is similar to 1D CNN but with 2D spatial operations. However, CNN architectures are not interpretable as their decisions are not understandable by human. Kolmogorov Arnold Networks (KANs) are the novel interpretable type of architectures. We investigate them in the Design Space to bring interpretability to unknown emitter detectors. The single KAN layer can be described as a MLP Layer, where activations are replaced with learnable functions and fixed weights on the edges. The learnable activation functions are often drawn from the k-order splines functions [de1978practical], which can be parameterized as linear combinations of basics - B-splines. These linear coefficients are then optimized during the training process. We implement a single-layer Efficient KAN version as:
@@
For CNNs and KANs we also evaluate SVD initialized modifications of those architectures. Let $F$ be a Feature Learning Model (CNN or KAN), which transforms the input data to $d$-dimensional features space. We initialize linear layer with SVD weights with the number of latent components $d$ and combine this linear layer with initial feature extractor $F$:
@@
where $L$ is the linear layer initialized with SVD weights. Under conditions of sparse and limited data like in 2D-Constellation data modality, this modification aims to provide additional knowledge of linear nature of the RF fingerprint (see Eq. 2) and boost the performance of complex self-supervised approach, which might degrade under above mentioned conditions [arora2017simple, fournier2019empirical]. This transformation does not move KAN model out of KAN architectures, because adding linear functions to cubic splines produces cubic splines.
@@ IV-D Machine Learning Approach
There are three common Machine Learning approaches used in Self-Supervised SEI: Deep Clustering, Auto Encoder and Contrastive Learning. In the current paper, we investigate the design dimension of Machine Learning Approach by evaluating mentioned above approaches.
@@ IV-D1 Deep Clustering
Deep Clustering [Caron_2018_ECCV] is a classical approach in unsupervised and self-supervised representation learning. Let $F$ denote the feature extractor mapping input $X$ to features $H=F(X)$. A classification head $M_{\text{class}}$ maps $H$ to $C$ predefined clusters. Training alternates between:
  • Clustering Step, which extracts features $H$ and clusters them with K-means into $C$ clusters, assigning pseudo-labels $p_{i}$.
  • Feature Learning Step, which updates $F$ and $M_{\text{class}}$ by minimizing cross-entropy between predictions and pseudo-labels.
@@ IV-D2 Auto Encoders
AEs [liu2021self] learn representations by reconstructing input tensors. The architecture is typically a U-net without skip connections, enforcing full encoding into the latent space. The loss function used is the mean square error (MSE) between the input $B$ and its reconstruction $B_{\text{rec}}$. In AE algorithm, learnable decoders are required. We use CNN with de-convolutions for CNN architecture and KAN layer for KAN architecture.
@@ IV-D3 Contrastive learning
We use common in SEI SimCLR [chen2020simple, hao2023contrastive] framework for contrastive learning, which works as following. For each sample $B_{i}$, two augmented views $B_{i}^{*},B_{i}^{**}$ are generated via random transformations. Features are extracted with $F$:
$H_{i}^{*}=F(B_{i}^{*}),\quad H_{i}^{**}=F(B_{i}^{**}).$ (18)
Using cosine similarity, the loss encourages positive pairs $(H_{i}^{*},H_{i}^{**})$ while contrasting against negatives:
$L=\sum_{i=1}^{N}\frac{e^{D(H_{i}^{*},H_{i}^{**})}}{\sum_{j\neq i}e^{D(H_{i}^{*},H_{j}^{*})}}+\sum_{i=1}^{N}\frac{e^{D(H_{i}^{*},H_{i}^{**})}}{\sum_{j\neq i}e^{D(H_{i}^{**},H_{j}^{**})}}.$ (19)
@@ IV-E Decision Making Module
Figure 5: Decision Making on a test sample. Three circles represent training clusters with probability densities equal to zero outside the circle. The red point represents a test sample, which in the current case is an outlier and predicted as unknown. $F1,F2$ are axes of demonstrative 2D features space.
There are two types of DMM in SEI: classification based and features space based. The first one is used for supervised tasks via classification head, and the second one employees different criteria in features space distribution and fits the self-supervised paradigm. We use clusters based approach as one of most widely used for unknown detection tasks and absolutely universal: it doesn’t depend on certain type of Features Learning Module or ML approach.
The clusters based DMM is design as following. After the training, we extract training features by trained features extractor $F$ and cluster them into $C$ clusters using K-means algorithm. During the inference, we extract features of test sample and detect the nearest cluster. Then, we compare the distance from the test sample to the nearest cluster center with the distances of train samples in this cluster. If a test sample lies in the right $\alpha\%$ quantile of the train distribution, it is marked as unknown. This concept is illustrated in Figure 5.
@@
If $s_{i}>\alpha$, then the test sample is labeled as unknown. Classification scores from Eq. (20) and nearest clusters indices are farther used for metrics computation.
Figure 6: WiSig and ORACLE datasets comparison. First columns shows I components and the second column represents Q components of the I/Q samples. Different colors refer to different signal traces from the datasets.
@@ V-A Datasets
As mentioned in Sect. IV-A, we evaluate the proposed SEI workflow under two scenarios with all emitters transmitting either the same or different messages. WiSig [hanna_wisig_2022] and ORACLE [sankhe2019oracle] datasets are designed for same and different messages scenarios, respectively, sharing similar indoor environment conditions.
@@ V-A1 WiSig
The WiSig dataset specifically includes signals captured from six distinct WiFi transmitting devices, recorded over a period of four consecutive days in indoor conditions. For each device, $1,000$ signals were recorded per day, resulting in a well-structured dataset designed to facilitate emitter identification tasks. Each signal trace consists of 256 consecutive I/Q samples, capturing a short time window of the emitted waveform. The upper line in Figure 6 shows I and Q components of two signal traces from the WiSig dataset; repeating patterns indicate that the messages of two signal traces are the same.
@@ V-A2 ORACLE
The ORACLE dataset was also collected in indoor conditions. It consists of 16 emitters each containing two long signals of length 4 millions recorded at the distance of 20 meters. Those signals are sliced into short traces of 256 I/Q samples each, similar as in the WiSig dataset. The number of samples per emitter is reduced to be the same as in the WiSig dataset, i.e. $4,000$. The bottom line in Figure 6 shows I and Q components of two signal traces from the ORACLE dataset; different patterns indicate that the messages of two signal traces are different.
@@ V-B Metrics
For evaluation in terms of clustering quality and detection performance we use three complementary metrics: the area under the Receiver Operating Characteristic Curve (ROC-AUC), Normalized Mutual Information (NMI) and F1-score. We calculate these metrics using predictions from DMM (see Sect. IV-E).
ROC-AUC metric measures the ability of the model to discriminate between classes across varying decision thresholds. A higher AUC value indicates better separability between known and unknown samples. Unlike F1, ROC-AUC provides a threshold-independent evaluation of detection quality. For ROC-AUC calculation, classification scores from DMM are used with $\{known(0),unknown(1)\}$ labels in the test dataset.
Normalized Mutual Information (NMI) metric quantifies the agreement between predicted cluster assignments and ground-truth labels. It measures how much information is shared between the predicted and true label distributions, where 0 indicates no mutual information and $1$ denotes a perfect clustering alignment. NMI is calculated between assigned clusters’ indices from DMM and emitters’ IDs.
F1-score metric represents the harmonic mean between precision and recall, defined as $F1=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$. It balances false positives and false negatives, making it an effective indicator of detection reliability in both closed- and open-set conditions. In our context, the F1-score is computed for the binary classification task of distinguishing between known and unknown samples. We get predictions via the procedure from DMM with $\alpha=5\%$.

### v2

@@ III-B Same Messages Scenario (SMS)
$\begin{split}&x_{i}(t)-x^{\prime}_{i}(t)\approx 0\\ &x_{i}(t)-x_{j}(t)=r_{emitter,i}(t)-r_{emitter,j}(t)\neq 0.\end{split}$ (4)
This makes the self-supervised task easier because of linearly separable classes.
III-C Different Messages Scenario (DMS)
In a more generic use case when we do not know the exact technology and protocol used, we may only have access to some random part of the received frame, which means that the clean signal $x_{symbols}$ is also unknown. The signal can now be expressed as:
@@ III-C Different Messages Scenario (DMS)
where $\hat{x}_{symbols}(t)$ is random. Thus, the variation within the emitter becomes the same as differences between emitters because the emitter specific part represents a small perturbation compared to the clean symbol as stated in Eq. 2 and expressed as follows:
@@ III-D Problem Statement
In this work, we formulate the problem of UED as a zero-shot learning task within a self-supervised learning setting. We develop an approach that extracts the RF fingerprint part from signals, as shown in Eq. 2, grouping identical emitters and separating different ones in the feature space. The objective is to learn a composite function
$\Phi:X_{\text{in}}\rightarrow L,\quad\text{where }L=\{\text{known}\>(0),\text{unknown}\>(1)\}$ (7)
@@
The composite function $\Phi(X_{\text{in}})$ from Eq. (7) can be decomposed into two modular sub-functions as per Eq. (8): the embedding function $F$, which transforms raw RF data into a projection (i.e. embedding) in a discriminative feature space, and the detection function $M$, which determines whether an embedded sample belongs to a known or an unknown emitter.
@@ III-D1 Embedding Function
Let $F$ be the embedding function trained in a self-supervised manner to project input signals into a latent space $F:X_{\text{in}}\rightarrow\mathbb{R}^{d}$, where $X_{\text{in}}$ denotes the input signal domain, and $\mathbb{R}^{d}$ is the latent feature space of dimension $d$. The output embeddings should cluster tightly for samples from the same emitter while remaining well-separated for different emitters. In other words, the latent space has to represent fingerprint components of signals from Eq. 2.
@@ III-D2 Detection Function
The detection function $M$ operates over the embedded space to determine whether a given sample is associated with a known or unknown emitter $M:\mathbb{R}^{d}\rightarrow L$.
IV Design of the UED workflow
To address the problem identified in Sec. III-D, we formalize the structured workflow for ML-based UED as depicted in Figure 3. The most important components of this workflow are: a) data modality, b) feature learning module (FLM), c) machine learning approach and d) decision making module. Data modality component transforms the input signal to the chosen representation: raw I/Q or 2D constellation. The FLM and machine learning approach represent the self-supervised learning block. The selected FLM: KAN or CNN is trained by one of the machine learning approaches: DC,CL, AE in a self-supervised manner. Principal component analysis (PCA) is a prominent baseline in unsupervised learning, but it does not fit within a feature learning model-ML approach scheme. Thus, we treat it as an individual feature-learning module with a specialized training procedure. The decision making module is the final stage of the ML workflow, where a classification decision is made based on extracted features. The other components include evaluating metrics and interpreting the FLM and DMM. While not being directly involved, they are essential for the performance measuring and analysis of the system. The following subsections provide a detailed design description of each component.
@@ IV-A Transmitting devices
In this study, devices transmit signals following two scenarios: the Same Messages Scenario as described in Sec. III-B and Different Messages Scenario as per Sec. III-C. The first scenario, in which all emitters transmit the same content all the time, is considered in the case where one has access to the preamble. The second scenario with different content is more realistic but also more challenging. The particular scenario affects the best design choices of Unknown Emitter self-supervised zero-shot detectors.
IV-B Data Modality
@@
In the receiver, the RF signal is discretized with the same time step: $x[i]=x(i\delta t)$, where $x[i]$ is a complex value denoting I and Q components. The way it is processed to be fed to the features learning module $F$ is referred to as a data modality, an important part of the design space. In this work, we investigate two types of data modalities: raw I/Q data and a 2D constellation. The raw I/Q is the convenient choice for OFDM data; each observation $x[i]=x[i]^{I}+jx[i]^{Q}\in\mathbb{C}^{N}$ is represented as a two-channel input: $X_{\text{in}}^{0i}=x^{I}[i],\quad X_{\text{in}}^{1i}=x^{Q}[i]$. In this study, we argue that the raw I/Q data modality produces poor quality under the realistic scenario, assuming different transmitting messages as described in Sec. III-C because of the prevalence of semantic information $x_{symbols}(t)$ over the fingerprint information $r_{emittter}(t)$ according to Eq. 5. To address this issue, one can transform the data in a time-invariant manner:
@@
where $T$ is a transformation operation over raw I/Q data, and $\sigma$ is a permutation operation over the time axis. The $T(X_{\text{in}})$ tensor does not contain any semantic information, because all messages have the same representation, assuming that transmitted signals are random. A straightforward way to provide the time-invariant property is to represent raw I/Q samples as a 2D constellation data. Firstly, the data is scaled to [1,1][-1,1] range:
@@
Secondly, the procedure counting I/Q observations over cells created by the grid on a 2D plane is applied as depicted in the Figure 4(a):
$\begin{split}T(X_{\text{in}})^{i,j}=\#\{k:\frac{X_{\text{in norm}}^{0k}+1}{2}\in[i\epsilon,(i+1)\epsilon],\\ \frac{X_{\text{in norm}}^{1k}+1}{2}\in[j\epsilon,(j+1)\epsilon]\}/N\\ \text{where }i,j\in[0,K-1],\epsilon=\frac{1}{K},\end{split},$ (11)
where $K$ is the grid size and $N$ is the number of I/Q observations. The resulting tensor looks like a heat map shown in Figure 4(b).
@@ IV-C Feature Learning Module
The FLM $F$ is a core part of self-supervised zero-shot unknown emitter detectors. Based on the available literature [liu2023overcoming, hao2023contrastive], Convolutional Neural Networks (CNNs) based architectures are employed for FLM s. To provide a more in-depth exploration of design principles of this component, we also study new, interpretable KAN based extractors and PCA baselines.
With respect to CNNs, we investigate 1D and 2D architectures for raw I/Q and 2D constellation data modalities, respectively. The CNN-1D two-channel network, adopted from [o2018over], consists of four blocks, each containing convolution, batch normalization, and max pooling operations. 2D CNN is similar to CNN-1D but with 2D spatial operations. However, CNN architectures are not interpretable, as their decisions are not understandable by humans.
As a novel, interpretable type of architecture, we investigate Kolmogorov-Arnold Networks (KAN s) [liu2024kan] in the design space to bring interpretability to unknown emitter detectors. The single KAN layer can be described as an MLP layer, where activations are replaced with learnable functions and fixed weights on the edges. The learnable activation functions are often drawn from the k-order spline functions [de1978practical], which can be parameterized as linear combinations of basic B-splines. These linear coefficients are then optimized during the training process. We develop a single-layer Efficient KAN version as:
@@
For CNNs and KAN s, we also investigate SVD initialized modifications of those architectures. Let $F$ be a FLM (CNN or KAN), which transforms the input data to $d$-dimensional features space. We initialize a linear layer with SVD weights with the number of latent components $d$ and combine this linear layer with the initial FLM $F$:
@@
where $L$ is the linear layer initialized with SVD weights. Under conditions of sparse and limited data, like in 2D constellation data modality, this modification aims to provide additional knowledge of the linear nature of the RF fingerprint (see Eq. 2) and boost the performance of a complex self-supervised approach, which might degrade under the conditions as mentioned earlier [arora2017simple, fournier2019empirical]. This transformation does not move the KAN model out of KAN architectures, because adding linear functions to cubic splines produces cubic splines.
We also investigate PCA - a prominent baseline in unsupervised learning. It builds a linear transformation that extracts low-dimensional, most informative features based on their contributions to total variation.
@@ IV-D Machine Learning Approach
According to Sec. II, there are three main ML approaches used in self-supervised SEI/UED: Deep Clustering, Auto Encoder, and Contrastive Learning. These approaches are used to train the FLM and are typically followed by a decision making module at inference time. In this paper, we investigate the design dimension of the ML approach by evaluating the performance trade-offs of the three selected approaches.
@@ IV-D1 Deep Clustering
DC [Caron_2018_ECCV] is an ML approach in self-supervised learning, which is employed in [krasnov2025novel]. The training procedure consists of two alternating stages. During the first stage, features are extracted from the training set using the current state of FLM and clustered using the K-means algorithm, assigning cluster numbers (pseudo-labels) to training samples. At the second stage FLM is updated using the classification objective with those pseudo-labels. The clustering results with label assignment are used solely for the FLM $F$’s training objective and are not reused thereafter for inference or decision making.
@@ IV-D2 Auto Encoders
AE [liu2021self] is a classical self-supervised ML approach, used in [huang2022deep]. It learns the FLM through introducing a decoder network and a reconstruction objective. The loss function used is the mean square error (MSE) between the input and its reconstruction. In the AE approach, learnable decoders are required. We use CNN with deconvolutions as the decoder for the CNN architecture and the KAN layer as the decoder for the KAN architecture.
@@ IV-D3 Contrastive learning
We investigate the SimCLR [chen2020simple] CL approach similar to [hao2023contrastive], while [sun2025few, liu2023overcoming, wu2023specific] also employ CL but in slightly different set-ups. For each sample, two augmented views are generated via random transformations. Views of the sample are treated as positive pairs, and others are treated as negative pairs. Features are extracted from augmented samples using FLM, and a cross-entropy-like loss computed with objective to classify positive and negative pairs.
Figure 5: Decision Making on a test sample. Three circles represent training clusters with probability densities equal to zero outside the circle. The red point represents a test sample, which, in this case, is an outlier and is predicted as unknown. $F1,F2$ are axes of demonstrative 2D features space.
@@ IV-E Decision Making Module
DMM is the final stage of ML based UED as depicted in Figure 3. It has to be independent of the choice of a specific FLM or ML approach. There are several widely used DMM s like K-means based [smiti2020critical], KNN based [hao2023contrastive], MLP-based [krasnov2025novel], etc. As we solve the SSL problem of outliers detection, we investigate the cluster-based DMM [smiti2020critical], which is a robust, universal, and interpretable method for outlier detection. After the training, we extract the training features using a trained FLM $F$ and cluster them into $C$ clusters with the K-means algorithm. During inference, we extract features from the test sample and identify the nearest cluster. Then, we compare the distance from the test sample to the nearest cluster center with the distances of the train samples in this cluster. If a test sample lies in the right $\alpha\%$ quantile of the train distribution, it is marked as unknown. This concept is illustrated in Figure 5.
@@
If $s_{i}>\alpha$, then the test sample is labeled as unknown. Classification scores from Eq. (15) and the indices of nearest clusters are further used for metric computation.
@@ V-A Datasets
As mentioned in Sec. IV-A, we leverage the ML-based UED workflow under two scenarios to study the design principles of zero-shot self-supervised detection. For the study we rely on the WiSig [hanna_wisig_2022] and ORACLE [sankhe2019oracle] datasets designed for the same and different messages scenarios, respectively, sharing similar indoor environment conditions.
Figure 6: WiSig and ORACLE datasets comparison. The first column shows I components, and the second column represents Q components of the I/Q samples. Different colors refer to different signal traces from the datasets.
@@ V-A1 WiSig
The WiSig dataset specifically includes signals captured from six distinct WiFi transmitting devices, recorded over a period of four consecutive days in indoor conditions. For each device, $1,000$ signals were recorded per day, resulting in a well-structured dataset designed to facilitate emitter identification tasks. Each signal trace consists of 256 consecutive I/Q samples, capturing a short time window of the emitted waveform. The upper line in Figure 6 shows the I and Q components of two signal traces from the WiSig dataset; the repeating patterns indicate that the messages of the two signal traces are the same.
@@ V-A2 ORACLE
The ORACLE dataset was also collected in indoor conditions. It consists of 16 emitters, each containing two long signals of 40 million I/Q observations recorded at a distance of 20 meters. Those signals are sliced into short traces of 256 I/Q samples each, similar to those in the WiSig dataset. The number of samples per emitter is reduced to match the WiSig dataset, i.e., $4,000$. The bottom line in Figure 6 shows I and Q components of two signal traces from the ORACLE dataset; different patterns indicate that the messages of two signal traces are different.
@@ V-B Metrics
For evaluation in terms of clustering quality and detection performance, we use three complementary metrics, i.e. the area under the Receiver Operating Characteristic curve (ROC-AUC), Normalized Mutual Information (NMI), and F1-score. We calculate these metrics using predictions from DMM described in Sec. IV-E.
The ROC-AUC metric measures the ability of the model to discriminate between classes across varying decision thresholds. A higher AUC value indicates better separability between known and unknown samples. Unlike F1, ROC-AUC provides a threshold-independent evaluation of detection quality. For ROC-AUC calculation, classification scores from DMM are used with $\{known(0),unknown(1)\}$ labels in the test dataset.
The NMI metric quantifies the agreement between predicted cluster assignments and ground-truth labels. It measures how much information is shared between the predicted and actual label distributions, where 0 indicates no mutual information and $1$ denotes a perfect clustering alignment. NMI is calculated between assigned clusters’ indices from DMM and emitters’ IDs.
F1-score metric represents the harmonic mean between precision and recall, defined as $F1=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$. It balances false positives and false negatives, making it an effective indicator of detection reliability in both closed- and open-set conditions. In our context, the F1-score is computed for the binary classification task of distinguishing between known and unknown samples. We obtain predictions via the procedure from DMM with $\alpha=5\%$. Along with performance metrics, the numbers of learnable parameters and floating point operations (FLOPs) are calculated for each FLM.